OpenStudy (anonymous):

lest common multiple of 42, 70?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

42=7*3*2 70=7*5*2 so LCM=7*3*5*2 =210

OpenStudy (anonymous):

2 is wrong